Rolling a Die If a die is rolled one time, find

these probabilities.

a. Getting a 2

b. Getting a number greater than 6 0

c. Getting an odd number

d. Getting a 4 or an odd number

e. Getting a number less than 7 1

f. Getting a number greater than or equal to 3

g. Getting a number greater than 2 and an even number


1
Expert's answer
2021-05-07T11:32:03-0400

S={x:x∈{1,2,3,4,5,6}}

a. the probability of getting a 2

P(x=2)=1/6

b. the probability of getting a number greater than 6

P(x>6)=0/6=0

c. the probability of getting an odd number

P(x=1)+P(x=3)+P(x=5)=1/6+1/6+1/6=3/6

d. the probability of getting a 4 or an odd number

P(x=4)+P(x=1)+P(x=3)+P(x=5)=1/6+1/6+1/6+1/6=4/6

e. the probability of getting a number less than 7

P(x<=7)=P(s)=1

f. the probability of getting a number greater than or equal to 3

P(x=3)+P(x=4)+P(x=5)+P(x=6)=1/6+1/6+1/6+1/6=4/6

g. the probability of getting a number greater than 2 and an even number

P(x=4)+P(x=6)=1/6+1/6=2/6